# Molecule Graph Networks with Many-body Equivariant Interactions

### v2

@@ 2.1 Equivariance
Recently, many works have leveraged group representation theory to design and study deep neural networks, aiming to incorporate more geometric information into deep learning models. This integration has enabled the deep learning architectures to effectively exploit the underlying symmetries and spatial relationships within data, further advancing their capability to model complex systems and interactions [30, 31, 32].
@@ 3.2 Characterizing representability with line graphs
which represents a combination of directional vectors $\vec{\mathbf{r}}_{\bullet\bullet}$ with tunable coefficients for each vector. Here, the $2$-body set $\{x,y\}$ represents a vertex in the line graph $\mathcal{L}[\mathcal{G}]$, corresponding to the edge $\{x,y\}$ in the graph $\mathcal{G}$. This calculation demonstrates that the use of line graphs allows the GNN framework to incorporate higher-order geometric interactions, such as angular relationships and edge dependencies, which are otherwise missed in traditional graph-based aggregation schemes.
@@
where $\vec{\mathbf{h}}_{i}$ denotes the directional vertex feature of node $i$. This formula provides a hierarchical aggregation framework that includes different levels of interaction relationships among atoms, namely: the atom level ($1$-body), bond level ($2$-body), and triplet level ($3$-body) aggregations. Fig. 4 illustrates this hierarchical aggregation within the proposed ENINet architecture. Further details on implementing this aggregation framework are provided in the Methods section (Section 5).
More generally, beyond the $3$-body interactions captured by line graph representations, higher-dimensional interactions, i.e., $N$-body interactions, can be iteratively modeled by considering higher-dimensional line graph structures. Formally, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with directional features $\mathcal{F}_{\mathcal{V}}=\{\vec{\mathbf{f}}_{v}\}_{v\in\mathcal{V}}$ and =$\mathcal{F}_{\mathcal{E}}=\{\vec{\mathbf{f}}_{e}\}_{e\in\mathcal{E}}$ on the vertices and edges, respectively, the aggregation process of vertex features on this graph is formalized through the following update rule:
@@
where $\mathrm{Agg}(\mathcal{G}^{(n)},\mathcal{F}_{\mathcal{V}^{(n)}},\mathcal{F}_{\mathcal{E}^{(n)}})$ updates $\mathcal{F}_{\mathcal{V}^{(n)}}$, and the updated vertex features are treated as edge features in the $(n-1)$-th level feature aggregation.
This process extends the hierarchical aggregation framework implemented in this work, incorporating progressively higher-order interaction relationships among the nodes. The framework naturally generalizes the aggregation to include $N$-body interactions, iteratively incorporating the information from higher-dimensional line graph structures. This iterative approach provides a flexible and powerful mechanism for modeling complex and intricate relationships in graph-structured data, such as molecular systems.
@@ 3.3 Quantum chemical properties
Table 1 compares the performance of ENINet with prior models on the QM9 test set. ENINet demonstrates superior accuracy across most properties, particularly for energy-related properties. This performance surpasses that of previous models employing invariant architectures and equivariant frameworks based on coordination vectors, which correspond to $l=1$ irreducible representations in the context of spherical harmonics.
@@
Molecule SchNet DimeNet PaiNN NequIP($l=1$) ET NewtonNet ENINet
@@
The total polarizability tensor is decomposed into two irreducible components for error measurement: the scalar ($\lambda=0$) $\alpha^{(0)}=\frac{\alpha_{11}+\alpha_{22}+\alpha_{33}}{\sqrt{3}}$ and the tensor ($\lambda=2$) $\boldsymbol{\alpha}^{(2)}=\sqrt{2}\left[\alpha_{12},\alpha_{23},\alpha_{13},\frac{2\alpha_{33}-\alpha_{11}-\alpha_{22}}{2\sqrt{3}},\frac{\alpha_{11}-\alpha_{22}}{2}\right]$ following [55].
Table 4 compares the performance on molecular polarizabilities. The reference $\boldsymbol{\alpha}$ values were obtained using LR-CCSD with the d-aug-cc-pVDZ basis set [59], which significantly reduces basis set incompleteness error at the double-$\zeta$ level and is proven capable of providing accurate and reliable predictions. The CCSD/DFT(B3LYP) error reveals a significant discrepancy between the two calculation methods. Notably, both AlphaML and ENINet exhibit lower errors than DFT when predicting CCSD values. ENINet outperforms AlphaML in predicting total values and tensorial components, although it slightly underperforms AlphaML in scalar predictions, for both CCSD and DFT tasks. Both models show higher RMSE values for DFT(B3LYP) calculations, as LR-CCSD with the d-aug-cc-pVDZ basis set serves as a more reliable reference for comparison.
@@ 4.2 Computational cost
Table 6 summarizes the computational costs of various high-order equivariant models when training on the Cv property of 133k small molecules from the QM9 dataset. For comparison, we set $l=2$ for the other models, although the most accurate results are typically achieved with even higher-order irreps. All models use a cutoff distance of $5\AA$ for graph construction and a training batch of 32. Training was performed on a single NVIDIA A100 40GB GPU. The comparison metrics include GPU memory usage, training time per epoch, and the number of trainable parameters. The introduction of three-body equivariant interactions in ENINet-T increases training time by 1.4 times and GPU memory usage by approximately 1.7 times. Despite these increases, ENINet-T remains both memory- and runtime-efficient compared to models that incorporate $l=2$ tensor symmetries.
@@ Conclusions
To enhance model expressivity while maintaining computational efficiency, we propose the use of equivariant coordination vectors in Cartesian space instead of high-order spherical harmonics, resulting in improvements in both memory and runtime efficiency. By leveraging only $l=1$ irreps, ENINet demonstrates superior performance across various benchmark datasets, outperforming traditional equivariant models in terms of predictive accuracy. Notably, ENINet achieves much lower errors even with smaller conformational training datasets, showcasing its ability to effectively capture complex energy relationships and conformational dependencies in molecular systems with limited data. This work presents an alternative framework that has the potential to enhance atomistic simulations with minimal additional computational cost.
@@ 5.1 Molecule representations
Similarly, the edge set is expressed as =$\mathcal{E}=\{(\mathbf{e}_{ij},\vec{\mathbf{e}}_{ij})\}^{1:N^{e}}$, where $\mathbf{e}_{ij}\in\mathbb{R}^{d\times 1}$ and $\vec{\mathbf{e}}_{ij}\in\mathbb{R}^{d\times 3}$ represent bond features derived from positional information between atom $i$ and $j$. The absolute positions of the atoms generate the relative position $\vec{\textbf{r}}_{ij}$ for each atom pair $(i,j)$ within the cutoff range $R_{c}$. A limitation of maximum number of neighbors $n_{\mathrm{max}}=32$ is set for each atom. The relative position is decomposed into two components: the distance $\|\vec{\textbf{r}}_{ij}\|$ and the direction $\vec{\textbf{r}}_{ij}/\|\vec{\textbf{r}}_{ij}\|$. The scalar bond feature $\mathbf{e}^{0}_{ij}$ is initialized by embedding the interatomic distance using a Gaussian basis expansion and multiplied by a cosine cutoff function $f_{\mathrm{cut}}$ to ensure gradient smoothness.
@@ 5.3 Equivariance analysis
In this work, we provide the mathematical foundations and theoretical analysis of the ENINet architecture in the supplementary information. By leveraging key concepts such as graph and line graph structures, tensor products, group representation theory, and equivariant/invariant functions, we rigorously define and analyze the ${\rm O}(3)$-equivariance of ENINet’s feature aggregation functions for 2-body, 3-body, and generalized $N$-body interactions. Incorporating line graphs and coordinate embeddings, we demonstrate how directional information and higher-dimensional interactions are seamlessly integrated to preserve equivariance. Through Theorems S1, S2, and S3, we establish the ${\rm O}(3)$-equivariance of the core aggregation functions, validating ENINet’s capability to handle molecular structure and property prediction tasks with geometric consistency.
@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Section A presents the mathematical foundations and analysis of the ENINet architecture. It introduces core concepts, including graph and line graph structures, tensor products, group representation theory, and equivariant/invariant functions, to rigorously establish the theoretical underpinnings of the ENINet model (Theorems S1, S2, and S3). These tools enable the formal definition and analysis of ${\rm O}(3)$-equivariant functions in feature aggregation, a key aspect of the proposed architecture.
@@
Section A.1 briefly introduces the concepts and notations of graphs and their line graphs utilized in this paper. Specifically, this work considers finite, undirected, simple graphs—i.e., finite undirected graphs with no self-loops or multiple edges—denoted by $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is a finite set of vertices, typically labeled as $1,2,\dots$, and $\mathcal{E}$ is the set of edges, represented as sets $\{i,j\}$ with distinct $i,j\in\mathcal{V}$.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected, simple graph. The set of neighbors of a vertex $v\in\mathcal{V}$ are defined as the set $\mathcal{N}(v)=\{u\in\mathcal{V}\mid\{u,v\}\in\mathcal{E}\}$. In particular, since $\mathcal{G}$ is a simple graph, $\mathcal{N}(v)$ does not include the vertex $v$ itself.
Beyond the connection relationships between vertices, the underlying graph of the ENINet architecture is also enriched with positional information for the vertices; in other words, it is a graph embedded in the three-dimensional Euclidean space $\mathbb{R}^{3}$. Specifically, the underlying graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is equipped with a family $\{\mathbf{r}_{i}\mid i\in\mathcal{V}\}$, representing the 3D coordinates of the vertices embedded in $\mathbb{R}^{3}$. In particular, the directional vector $\vec{\textbf{r}}_{ij}$ from vertex $i$ to vertex $j$ is typically defined as the subtraction $\textbf{r}_{j}-\textbf{r}_{i}\in\mathbb{R}^{3}$, which encodes the pairwise directional relationships between the embedded vertices.
@@
To leverage higher-dimensional interactions between vertices, the line graph is utilized in the proposed ENINet architecture. For a more comprehensive introduction to line graphs, refer to [62, 63]. Intuitively, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the associated line graph $\mathcal{L}[\mathcal{G}]$ is constructed by treating $\mathcal{E}$ as its vertex set, where two vertices in $\mathcal{L}[\mathcal{G}]$ are connected by an edge if the corresponding edges in $\mathcal{G}$ share a common endpoint.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected, simple graph. The line graph of $\mathcal{G}$, denoted as $\mathcal{L}[\mathcal{G}]$, is a graph that is defined as follows:
  • the vertex set of $\mathcal{L}[\mathcal{G}]$ is defined by $\mathcal{E}$;
  • two distinct vertex $\{i,j\},\{k,l\}\in\mathcal{E}$ is connected by the edge $\{\{i,j\},\{k,l\}\}$ if $\{i,j\}\cap\{k,l\}\neq\emptyset$.
Then, the line graph $\mathcal{L}[\mathcal{G}]$ is an undirected, simple graph.
Furthermore, based on the line graph construction, the “higher-dimensional” line graph can be defined iteratively [64]. In this work, we define $\mathcal{G}^{(0)}=\mathcal{G}$, $\mathcal{G}^{(1)}=\mathcal{L}[\mathcal{G}]$, and $\mathcal{G}^{(n+1)}=\mathcal{L}[\mathcal{G}^{(n)}]$ for $n\geq 0$, where $\mathcal{V}^{(n)}$ and $\mathcal{E}^{(n)}=\mathcal{V}^{(n+1)}$ denote the vertex and edge sets of $\mathcal{G}^{(n)}$, respectively. In other words, a sequence of finite, undirected, simple graphs is iteratively constructed and utilized in the ENINet architecture:
$\mathcal{G}^{(0)}=(\mathcal{V}^{(0)},\mathcal{E}^{(0)}),\ \ \mathcal{G}^{(1)}=(\mathcal{V}^{(1)},\mathcal{E}^{(1)}),\ \ \mathcal{G}^{(2)}=(\mathcal{V}^{(2)},\mathcal{E}^{(2)}),\ \ \dots,$
where $\mathcal{G}^{(0)}:=\mathcal{G}=(\mathcal{V},\mathcal{E})$. This iterative construction allows us to consider “higher-dimensional” interactions between atoms within a group of atoms.
Furthermore, with the inclusion of vertex coordinates in the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, position information can also be encoded into the line graph structure. Specifically, let $\{\mathbf{r}_{i}\mid i\in\mathcal{V}\}$ represent the coordinates of the vertices in $\mathcal{G}$. The position information of an edge $\{i,j\}$ can be defined as a linear combination of $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$, namely $\alpha\mathbf{r}_{i}+\beta\mathbf{r}_{j}$, for some real numbers $\alpha,\beta\in\mathbb{R}$. Similarly, for an edge in the line graph $\mathcal{L}[\mathcal{G}]$, with "endpoints" $\{i,j\}$ and $\{i,k\}$, the directional information can be defined in the same manner as in the original graph. In particular, this directional information is a linear combination of $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected, simple graph, and let $\nu^{(0)}:\mathcal{V}\rightarrow\mathbb{R}^{3}$ be an embedding function that assigns each vertex to its 3D coordinate. Let $\mathcal{G}^{(1)}=(\mathcal{V}^{(1)},\mathcal{E}^{(1)})$ denote the line graph of $\mathcal{G}$. The coordinate embedding of vertices in $\mathcal{G}^{(1)}$ is defined as a function $\nu^{(1)}:\mathcal{V}^{(1)}\rightarrow\mathbb{R}^{3}$, given by
@@
for some real numbers $\alpha_{i},\alpha_{j}\in\mathbb{R}$. In other words, each edge $\{i,j\}\in\mathcal{V}^{(1)}$ is encoded as a linear combination of the 3D coordinate vectors of its endpoints $i$ and $j$, as defined by $\nu^{(0)}$.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a finite, undirected, simple graph, and let $\nu^{(0)}:\mathcal{V}\rightarrow\mathbb{R}^{3}$ be an embedding function. Let $\mathcal{G}^{(1)}=(\mathcal{V}^{(1)},\mathcal{E}^{(1)})$ be the line graph of $\mathcal{G}$, and let $\nu^{(1)}:\mathcal{V}^{(1)}\rightarrow\mathbb{R}^{3}$ be the induced embedding function. If $Q\in\mathbb{R}^{3\times 3}$ is a $3\times 3$ matrix, and $\eta^{(0)}:\mathcal{V}\rightarrow\mathbb{R}^{3}$ is defined by $i\mapsto Q\cdot\nu^{(0)}(i)$, then the induced embedding $\eta^{(1)}:\mathcal{V}^{(1)}\rightarrow\mathbb{R}^{3}$ is exact the function $\{i,j\}\mapsto Q\cdot\nu^{(1)}(\{i,j\})$.
@@
By the definition of the embedding function $\eta^{(1)}$ of the line graph $\mathcal{G}^{(1)}=(\mathcal{V}^{(1)},\mathcal{E}^{(1)})$,
@@
Let $V$ and $W$ be $\mathbb{R}$-vector spaces with bases $\beta$ and $\gamma$, respectively. The tensor product of $V$ and $W$, denoted by $V\otimes W$, is the vector space generated by the family $\{(v,w)\mid v\in\beta,w\in\gamma\}$ as its basis. In the tensor product, an element $(v,w)$ with $v\in V$ and $w\in W$ is denoted by $v\otimes w$.
@@
The following example illustrates the decomposition of a tensor product of two (column) vectors by the standard basis of the tensor product of the underlying vector spaces. Specifically, for the vectors
@@
the associated tensor product of the vectors can be decomposed as the form
@@
of a linear combination of the standard basis of $\mathbb{R}^{4}\otimes\mathbb{R}^{3}$. By reshaping the coefficients, this tensor product can be also represented as the $4\times 3$ matrix

### v3

@@ 2.1 Equivariance
Recently, many works have leveraged group representation theory to design and study deep neural networks, aiming to incorporate more geometric information into deep learning models. This integration has enabled the deep learning architectures to effectively exploit the underlying symmetries and spatial relationships within data, further advancing their capability to model complex systems and interactions [30, 31, 32]. See the Supporting Information for a more detailed discussion of equivariance and group representation theory.
@@ 3.2 Characterizing representability with line graphs
which represents a combination of directional vectors $\vec{\mathbf{r}}_{\bullet\bullet}$ with tunable coefficients for each vector. Here, the 222-body set $\{x,y\}$ represents a vertex in the line graph $\mathcal{L}[\mathcal{G}]$, corresponding to the edge $\{x,y\}$ in the graph $\mathcal{G}$. This calculation demonstrates that the use of line graphs allows the GNN framework to incorporate higher-order geometric interactions, such as angular relationships and edge dependencies, which are otherwise missed in traditional graph-based aggregation schemes.
@@
where $\vec{\mathbf{h}}_{i}$ denotes the directional vertex feature of node $i$. This formula provides a hierarchical aggregation framework that includes different levels of interaction relationships among atoms, namely: the atom level (111-body), bond level (222-body), and triplet level (333-body) aggregations. Fig. 5 illustrates this hierarchical aggregation within the proposed ENINet architecture. Further details on implementing this aggregation framework are provided in the Methods section (Section 6).
More generally, beyond the 333-body interactions captured by line graph representations, higher-dimensional interactions, i.e., $N$-body interactions, can be iteratively modeled by considering higher-dimensional line graph structures. Formally, given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with directional features $\mathcal{F}_{\mathcal{V}}=\{\vec{\mathbf{f}}_{v}\}_{v\in\mathcal{V}}$ and =$\mathcal{F}_{\mathcal{E}}=\{\vec{\mathbf{f}}_{e}\}_{e\in\mathcal{E}}$ on the vertices and edges, respectively, the aggregation process of vertex features on this graph is formalized through the following update rule:
@@
[…]
This process extends the hierarchical aggregation framework implemented in this work, incorporating progressively higher-order interaction relationships among the nodes. The framework naturally generalizes the aggregation to include $N$-body interactions, iteratively incorporating the information from higher-dimensional line graph structures. This iterative approach provides a flexible and powerful mechanism for modeling complex and intricate relationships in graph-structured data, such as molecular systems. An introduction to the core concepts of graphs, line graphs, and tensor products is provided in the Supporting Information.
@@ 3.3 Quantum chemical properties
Table 1 compares the performance of ENINet with prior models on the QM9 test set. ENINet demonstrates superior accuracy across most properties, particularly for energy-related properties. This performance surpasses that of previous models employing invariant architectures and equivariant frameworks based on coordination vectors, which correspond to $l=1$ irreducible representations in the context of spherical harmonics. Although our model outperforms existing methods on QM9 across most properties, certain cases show limited gains, consistent with empirical evidence suggesting that higher-order interactions provide marginal benefit for small molecules. [52]. Their relatively simple interaction patterns are often sufficiently captured by simpler models, leading to a performance plateau where additional complexity yields diminishing returns. To further evaluate this, we conducted additional experiments across a broader range of tasks.
@@
[…]
@@
Solvated molecular force fields
Figure 4: (a) Comparison of ENINet, MACE ($l=1$), and MACE ($l=2$) in terms of force RMSE, energy MAE, GPU memory usage (MB), and inference time across five randomly selected test molecules with multiple conformers. (b) 3D structures of the five selected test molecules, shown in order of increasing molecular size.
Understanding solvent effects and collective dispersion interactions remains a critical challenge in computational drug design, as both strongly influence the conformational stability and quantum-mechanical properties of drug-like molecules. We evaluate the performance and scalability of ENINet in comparison with MACE models using $l=1$ and $l=2$ equivariant features on the Aquamarine (AQM) dataset [55]. This dataset comprises 59,786 conformers (both low- and high-energy) of 1,653 larger molecules containing up to 98 atoms. For solvated molecules, calculations were performed at the PBE0 level with many-body dispersion (MBD) corrections supplemented by the modified Poisson-Boltzmann (MPB) implicit solvent model. Energies and forces explicitly account for many-body dispersion effects in this task. All models use a cutoff radius of 4 Åto construct molecular graphs. The number of parameters is 7.9 million for ENINet and 1.2 million for MACE.
To assess scaling behavior, we randomly selected five molecules from the AQM dataset with increasing atom counts (14, 31, 50, 70, and 90 atoms), retaining all their conformers as test data. The models were trained on the conformers of the remaining molecules and then evaluated on the selected test conformers. The results, summarized in Fig. 4 (a), show that ENINet outperforms MACE ($l=1$) on forces for 4 out of 5 molecules and on energies for 3 out of 5. Compared to MACE ($l=2$), ENINet slightly underperforms, with lower errors observed in only 2 molecules for both forces and energies. This is likely because ENINet models higher-order interactions using simpler Cartesian vectors, which may lack sufficient expressivity to fully capture complex molecular interactions.
GPU memory usage indicates that ENINet’s resource demand falls between MACE ($l=1$) and MACE ($l=2$). Although ENINet employs $l=1$ equivariant features, its deeper convolutional layers and higher parameter count increase memory consumption relative to MACE ($l=1$), while the $l=2$ model incurs substantially higher computational load due to more complex higher-order tensor operations. Regarding inference time, ENINet is the fastest among the three models. This advantage is attributed to the computational expense of tensor product operations in MACE, particularly for higher-body and higher-order equivariant features. Finally, Fig. 4 (b) presents 3D visualizations of one representative conformer from each of the five test molecules, illustrating the diversity in molecular size and complexity.
@@
The total polarizability tensor is decomposed into two irreducible components for error measurement: the scalar ($\lambda=0$) $\alpha^{(0)}$ and the tensor ($\lambda=2$) $\boldsymbol{\alpha}^{(2)}$ following [57], where
$\alpha^{(0)}$ $=\frac{\alpha_{11}+\alpha_{22}+\alpha_{33}}{\sqrt{3}}$ (18)
$\boldsymbol{\alpha}^{(2)}$ $=\sqrt{2}\left[\alpha_{12},\alpha_{23},\alpha_{13},\frac{2\alpha_{33}-\alpha_{11}-\alpha_{22}}{2\sqrt{3}},\frac{\alpha_{11}-\alpha_{22}}{2}\right]$ (19)
[…]
@@ 4.2 Computational cost
Table 6 summarizes the computational costs of various high-order equivariant models when training on the Cv property of 133k small molecules from the QM9 dataset. For comparison, we set $l=2$ for the other models, although the most accurate results are typically achieved with even higher-order irreps. All models use a cutoff distance of 5 Åfor graph construction and a training batch of 32. Training was performed on a single NVIDIA A100 40GB GPU. The comparison metrics include GPU memory usage, training time per epoch, and the number of trainable parameters. The introduction of three-body equivariant interactions in ENINet-T increases training time by 1.4 times and GPU memory usage by approximately 1.7 times. Despite these increases, ENINet-T remains both memory- and runtime-efficient compared to models that incorporate $l=2$ tensor symmetries.
@@ 5 Conclusions
To enhance model expressivity while maintaining computational efficiency, we propose using equivariant coordination vectors in Cartesian space as an alternative to high-order spherical harmonics. This approach offers a different trade-off between improved accuracy, memory usage, runtime efficiency, and the complexity of higher-order tensor representations. By leveraging only $l=1$ irreps, ENINet demonstrates superior performance across various benchmark datasets, outperforming traditional equivariant models in terms of predictive accuracy. Notably, ENINet achieves much lower errors even with smaller conformational training datasets, showcasing its ability to effectively capture complex energy relationships and conformational dependencies in molecular systems with limited data. This work presents an alternative framework that has the potential to enhance atomistic simulations with minimal additional computational cost.
@@ 6.1 Molecule representations
[…]
@@ 6.3 Equivariance analysis
In this work, we provide the full mathematical foundations and theoretical analysis of the ENINet architecture in the supporting information. By leveraging key concepts such as graph and line graph structures, tensor products, group representation theory, and equivariant/invariant functions, we rigorously define and analyze the ${\rm O}(3)$-equivariance of ENINet’s feature aggregation functions for 2-body, 3-body, and generalized $N$-body interactions. Incorporating line graphs and coordinate embeddings, we demonstrate how directional information and higher-dimensional interactions are seamlessly integrated to preserve equivariance. Through Theorems S1, S2, and S3, we establish the ${\rm O}(3)$-equivariance of the core aggregation functions, validating ENINet’s capability to handle molecular structure and property prediction tasks with geometric consistency.
@@
Supporting information
The mathematical foundations underlying the design of the proposed architecture, such as line graphs, tensor products, group representation theory, and neural network equivariance, are detailed in the Supporting Information. Furthermore, the theoretical analysis of the ENINet architecture is presented in Theorems S1 to S3 within the same section.
@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Section A presents the mathematical foundations and analysis of the ENINet architecture. It introduces core concepts, including graph and line graph structures, tensor products, group representation theory, and equivariant/invariant functions, to rigorously establish the theoretical underpinnings of the ENINet model (Theorems S1, S2, and S3). These tools enable the formal definition and analysis of $\operatorname{O}(3)$-equivariant functions in feature aggregation, a key aspect of the proposed architecture.
@@
Section A.1 briefly introduces the concepts and notations of graphs and their line graphs utilized in this paper. Specifically, this work considers finite, undirected, simple graphs—i.e., finite undirected graphs with no self-loops or multiple edges—denoted by $\mathscr{G}=(\mathscr{V},\mathscr{E})$, where $\mathscr{V}$ is a finite set of vertices, labeled by integers 1,2,1,2$\dots$1 , 2 , … from the set $\mathbb{N}$ of natural numbers, and $\mathscr{E}$ is the set of edges, represented as sets $\{i,j\}$ with distinct $i,j\in\mathscr{V}$.
@@
Let $\mathscr{G}=(\mathscr{V},\mathscr{E})$ be an undirected, simple graph. The set of neighbors of a vertex $v\in\mathscr{V}$ are defined as the set $\mathscr{N}(v)=\{u\in\mathscr{V}\mid\{u,v\}\in\mathscr{E}\}$. In particular, since $\mathscr{G}$ is a simple graph, $\mathscr{N}(v)$ does not include the vertex $v$ itself.
Beyond the connection relationships between vertices, the underlying graph of the ENINet architecture is also enriched with positional information for the vertices; in other words, it is a graph embedded in the three-dimensional Euclidean space $\mathbb{R}^{3}$. Specifically, the underlying graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ is equipped with a family $\{\mathbf{r}_{i}\mid i\in\mathscr{V}\}$, representing the 3D coordinates of the vertices embedded in $\mathbb{R}^{3}$. In particular, the directional vector $\vec{\textbf{r}}_{ij}$ from vertex $i$ to vertex $j$ is typically defined as the subtraction $\textbf{r}_{j}-\textbf{r}_{i}\in\mathbb{R}^{3}$, which encodes the pairwise directional relationships between the embedded vertices.
@@
To leverage higher-dimensional interactions between vertices, the line graph is utilized in the proposed ENINet architecture. For a more comprehensive introduction to line (di)graphs, refer to [64, 65]. Intuitively, given a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$, the associated line graph $\mathscr{L}[\mathscr{G}]$ is constructed by treating $\mathscr{E}$ as its vertex set, where two vertices in $\mathscr{L}[\mathscr{G}]$ are connected by an edge if the corresponding edges in $\mathscr{G}$ share a common endpoint.
@@
Let $\mathscr{G}=(\mathscr{V},\mathscr{E})$ be an undirected, simple graph. The line graph of $\mathscr{G}$, denoted as $\mathscr{L}[\mathscr{G}]$, is a graph that is defined as follows:
  • the vertex set of $\mathscr{L}[\mathscr{G}]$ is defined by $\mathscr{E}$;
  • two distinct vertices $\{i,j\},\{k,l\}\in\mathscr{E}$ in the graph $\mathscr{L}[\mathscr{G}]$ is connected by the edge $\{\{i,j\},\{k,l\}\}$ if $\{i,j\}\cap\{k,l\}\neq\emptyset$.
In particular, the line graph $\mathscr{L}[\mathscr{G}]$ is an undirected, simple graph.
Furthermore, based on the line graph construction, the “higher-dimensional” line graph can be defined iteratively [66]. In this work, we define $\mathscr{G}^{(0)}=\mathscr{G}$, $\mathscr{G}^{(1)}=\mathscr{L}[\mathscr{G}]$, and $\mathscr{G}^{(n+1)}=\mathscr{L}[\mathscr{G}^{(n)}]$ for $n\geq 0$, where $\mathscr{V}^{(n)}$ and $\mathscr{E}^{(n)}=\mathscr{V}^{(n+1)}$ denote the vertex and edge sets of $\mathscr{G}^{(n)}$, respectively. In other words, a sequence of finite, undirected, simple graphs is iteratively constructed and utilized in the ENINet architecture:
$\mathscr{G}^{(0)}=(\mathscr{V}^{(0)},\mathscr{E}^{(0)}),\ \ \mathscr{G}^{(1)}=(\mathscr{V}^{(1)},\mathscr{E}^{(1)}),\ \ \mathscr{G}^{(2)}=(\mathscr{V}^{(2)},\mathscr{E}^{(2)}),\ \ \dots,$
where $\mathscr{G}^{(0)}:=\mathscr{G}=(\mathscr{V},\mathscr{E})$. This iterative construction allows us to consider “higher-dimensional” interactions between atoms within a group of atoms.
Furthermore, with the inclusion of vertex coordinates in the graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$, position information can also be encoded into the line graph structure. Specifically, let $\{\mathbf{r}_{i}\mid i\in\mathscr{V}\}$ represent the coordinates of the vertices in $\mathscr{G}$. The position information of an edge $\{i,j\}$ can be defined as a linear combination of $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$, namely $\alpha\mathbf{r}_{i}+\beta\mathbf{r}_{j}$, for some real numbers $\alpha,\beta\in\mathbb{R}$. Similarly, for an edge in the line graph $\mathscr{L}[\mathscr{G}]$, with "endpoints" $\{i,j\}$ and $\{i,k\}$, the directional information can be defined in the same manner as in the original graph. In particular, this directional information is a linear combination of $\mathbf{r}_{i}$ and $\mathbf{r}_{j}$.
@@
Let $\mathscr{G}=(\mathscr{V},\mathscr{E})$ be a finite, undirected, simple graph, and let $\nu^{(0)}:\mathscr{V}\rightarrow\mathbb{R}^{3}$ be an embedding function that assigns to each vertex a point in $\mathbb{R}^{3}$, representing its spatial coordinate. Let $\mathscr{G}^{(1)}=(\mathscr{V}^{(1)},\mathscr{E}^{(1)})$ denote the line graph of $\mathscr{G}$. The coordinate embedding of vertices in $\mathscr{G}^{(1)}$ is defined as a function $\nu^{(1)}:\mathscr{V}^{(1)}\rightarrow\mathbb{R}^{3}$, given by
@@
for some real numbers $\alpha_{i},\alpha_{j}\in\mathbb{R}$. In other words, each edge $\{i,j\}\in\mathscr{V}^{(1)}$ is encoded as a linear combination of the 3D coordinate vectors of its endpoints $i$ and $j$, as defined by $\nu^{(0)}$.
@@
Let $\mathscr{G}=(\mathscr{V},\mathscr{E})$ be a finite, undirected, simple graph, and let $\nu^{(0)}:\mathscr{V}\rightarrow\mathbb{R}^{3}$ be an embedding function. Let $\mathscr{G}^{(1)}=(\mathscr{V}^{(1)},\mathscr{E}^{(1)})$ be the line graph of $\mathscr{G}$, and let $\nu^{(1)}:\mathscr{V}^{(1)}\rightarrow\mathbb{R}^{3}$ be the induced embedding function. If $Q\in\mathbb{R}^{3\times 3}$ is a $\times 3$ 3 matrix, and $\eta^{(0)}:\mathscr{V}\rightarrow\mathbb{R}^{3}$ is defined by $i\mapsto Q\cdot\nu^{(0)}(i)$, then the induced embedding $\eta^{(1)}:\mathscr{V}^{(1)}\rightarrow\mathbb{R}^{3}$ is exact the function $\{i,j\}\mapsto Q\cdot\nu^{(1)}(\{i,j\})$.
@@
By the definition of the embedding function $\eta^{(1)}$ of the line graph $\mathscr{G}^{(1)}=(\mathscr{V}^{(1)},\mathscr{E}^{(1)})$,
@@
[…]
@@
The following example illustrates the decomposition of a tensor product of two (column) vectors by the standard bases of the tensor product of the underlying vector spaces. Specifically, for the vectors
@@
the associated tensor product of the vectors can be decomposed in the form
@@
of a linear combination of the standard basis of $\mathbb{R}^{4}\otimes\mathbb{R}^{3}$. By reshaping the coefficients, this tensor product can be also represented as the 4×34$\times$ 34 × 3 matrix